OpenStudy (anonymous):

will give medal. please help. Ella obtains a $750,000 balloon mortgage for an investment property. The mortgage has terms of 15/3 at an interest rate of 5.5%. The monthly payment is $6,128.13. How much is the remaining balance after 3 years rounded to the nearest dollar?

OpenStudy (anonymous):

@jim_thompson5910 @AccidentalAiChan

OpenStudy (anonymous):

@isag can u help me?

jimthompson5910 (jim_thompson5910):

According to the mortgage calculator from http://www.mlcalc.com/ I'm getting $648,106.64

jimthompson5910 (jim_thompson5910):

then round that to the nearest dollar to get $648,107

OpenStudy (anonymous):

that's not it :(

jimthompson5910 (jim_thompson5910):

dang...hmm, what are your answer choices?

jimthompson5910 (jim_thompson5910):

also, does it give you a formula to work with? It's possible they're altering the formula a bit

jimthompson5910 (jim_thompson5910):

According to this calculator http://tcalc.timevalue.com/all-financial-calculators/mortgage-calculators/balloon-mortgage-calculator.aspx it's saying that the balloon payment is $651,076.97 which rounds to $651,077

jimthompson5910 (jim_thompson5910):

So maybe that's what they want?

OpenStudy (anonymous):

$529,387 $644,949 $704,391 $600,000

OpenStudy (anonymous):

no, no formula. idk it

jimthompson5910 (jim_thompson5910):

ok one sec

jimthompson5910 (jim_thompson5910):

ok now I'm turning to the formula found here http://www.financeformulas.net/Loan_Balloon_Balance.html

jimthompson5910 (jim_thompson5910):

And I get... FV = PV(1+r)^n - P*((1+r)^n-1)/r FV = 750000(1+0.055/12)^(12*3) - 6128.13*((1+0.055/12)^(12*3)-1)/(0.055/12) FV = 644,948.841110999 FV = 644,949 ... rounding to the nearest dollar

jimthompson5910 (jim_thompson5910):

FV = future value (ie the remaining balance at some point in time) PV = present value (which is the original balance...aka the total loan amount) r = interest rate per period (per month in this case) n = number of payments P = monthly payment

jimthompson5910 (jim_thompson5910):

So in this case FV = unknown for now PV = 750,000 r = 0.055/12 n = 12*3 P = 6,128.13

OpenStudy (anonymous):

thank u so much !! can u help me with 2 more please

OpenStudy (anonymous):

Brandon has purchased a home for $352,000. He made a 15% down payment and financed the remaining amount. The intangible tax is 0.4%. Which of the following is the total amount of the intangible tax?

jimthompson5910 (jim_thompson5910):

The definition of intangible tax is a bit vague, but it sounds like we're just taking 0.4% of the current value of the home.

OpenStudy (anonymous):

yeah it is vague. they should explain better

jimthompson5910 (jim_thompson5910):

if so, then 352,000*(0.4/100) = $1,408 is the intangible tax

OpenStudy (anonymous):

thank u !!! An $800,000 balloon mortgage has terms of 30/5 with a 5% interest rate. What is the monthly payment rounded to the nearest dollar?

jimthompson5910 (jim_thompson5910):

using this calculator again http://tcalc.timevalue.com/all-financial-calculators/mortgage-calculators/balloon-mortgage-calculator.aspx it's saying: Payment amount $4,294.57

OpenStudy (anonymous):

Ridley obtains a 20-year 6.3% fixed rate mortgage for $323,500. The monthly payments are about $2,374. What is the total amount of interest for Ridleys mortgage?

OpenStudy (anonymous):

thanks so much !!

OpenStudy (anonymous):

that online calculator u are using.. I can never figure it out

jimthompson5910 (jim_thompson5910):

which one? I'm using a few calculators

jimthompson5910 (jim_thompson5910):

this is what I did for the last problem I posted

OpenStudy (anonymous):

oh ok perfect.. thanks ! I see

jimthompson5910 (jim_thompson5910):

notice how the down payment is 0% (ie we have NOT put any down payment down and we're financing all of the $800,000)

OpenStudy (anonymous):

A mortgage is in the amount of $300,000. The origination fee is 0.7%, the intangible tax is 0.3%, and there are 2.3 discount points. Which of the following is the total cost of the origination fee, intangible tax, and discount points? I have this one.. im going to try with that calculator and u tell me if I get it right

jimthompson5910 (jim_thompson5910):

and I'm sure you know this already, but when they say "30/5 balloon mortgage" they mean "treat it like a 30 year mortgage BUT the balloon payment happens at the start of the 5th year"

OpenStudy (anonymous):

oh ok.. didn't know that

jimthompson5910 (jim_thompson5910):

yeah it's a bit strange, but you pay normal monthly payments for the first 4 years, then you pay the balloon payment (which is the huge remaining balance) at the start of year 5

OpenStudy (anonymous):

I just don't know where to put 0.7% and 0.3 and 2.3 in the calculator

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

The bank offers Manfred a fixed rate loan with no points and a fixed rate loan with points. The difference in monthly payments between the loan with points and the loan without points is $38.42. The cost of the points is $2,300. What number is closest to the break-even point rounded to the nearest month? I have this one too. I cant wait to be done

jimthompson5910 (jim_thompson5910):

ok I did not know this, but apparently a discount point lowers the interest rate and they usually cost 1% of the loan. Unfortunately this isn't a hard set rule, but we'll go with it http://www.investopedia.com/terms/d/discountpoints.asp

jimthompson5910 (jim_thompson5910):

So if you purchase 2.3 discount points, then it will cost 2.3*1 = 2.3% of the loan

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Add up the percentage points: 0.3+0.7+2.3 = 3.3 Then take 3.3% of the amount loaned (3.3/100)*300,000 = 9,900 So it seems that the total cost of the origination fee, intangible tax, and discount points is $9,900

jimthompson5910 (jim_thompson5910):

This amount may be added to the balance or the amount you have to pay up front.

OpenStudy (anonymous):

okay. I see

OpenStudy (anonymous):

thank u !!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

do u know this one. The bank offers Manfred a fixed rate loan with no points and a fixed rate loan with points. The difference in monthly payments between the loan with points and the loan without points is $38.42. The cost of the points is $2,300. What number is closest to the break-even point rounded to the nearest month?

OpenStudy (anonymous):

that's my last one

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

let p1 be the payment made with no points let p2 be the payment made with points since the difference in the payments is $38.42, we know that p1 - p2 = 38.42 Now let y be the total cost so far after x months For the loan without points, we can say y = p1*x For the loan with points, we can say y = p2*x + 2300 ... note: remember to add on the cost of the points now we have this system y = p1*x y = p2*x + 2300 and subtract to get y = p1*x y = p2*x + 2300 ---------------- 0 = (p1-p2)*x - 2300 so we have 0 = (p1-p2)*x - 2300 0 = 38.42*x - 2300 2300 = 38.42*x 2300/38.42 = x 59.8646538261322 = x x = 59.8646538261322 So that rounds to 60 months (60/12 = 5 years)

OpenStudy (anonymous):

thank u !! I just submitted it. I got 26 out of 40 points. some we did were wrong :(

jimthompson5910 (jim_thompson5910):

so the last one wasn't 60 months?

jimthompson5910 (jim_thompson5910):

does it say which ones were wrong?